# Imitation from Diverse Behaviors: Wasserstein Quality Diversity Imitation Learning with Single-Step Archive Exploration

### v1

@@ 2.5 Encouraging exploration in RL and IL
Exploration bonuses have a long history in bandits and reinforcement learning as part of the “optimism in face of uncertainty” principle, and include but are not limited to visitation count based approaches. In bandits, upper confidence bound (UCB, e.g. [3]) computes a bonus based on $\sqrt{\frac{2\ln(t)}{n(a_{t})}}$ for action $a_{t}$ at time $t$. In tabular reinforcement learning, UCRL techniques (e.g. [24]) provide upper confidence bounds to the reward based on a bonus proportional to $\sqrt{\frac{1}{n(s_{t},a_{t})}}$. R-MAX [8] and E3 [25] explicitly distinguish between known and unknown states, where in unknown states R-MAX assigns the maximal reward of the MDP and E3 performs balanced wandering (i.e. taking the action with lowest visitation count for the current state). Related concepts have also been applied in deep reinforcement learning [42, 22, 29, 40]. Related to exploration bonuses, recent techniques in imitation learning also considered curiosity-based exploration [32, 9, 49], which relates to discrepancy between the transition model’s prediction and the true transition. For instance, GIRIL [49] computes rewards offline by pretraining a reward model using a conditional VAE [41].
@@
Our measure bonus differs from these approaches in that it is designed to explore the behavior space rather than the (state-)action space. Further, contrasting to UCRL type bonuses, due to dividing by the proportion of visits, rather than the count, our measure bonus does not shrink to zero such that underexplored behaviors will continue to receive a bonus. This is somewhat comparable in aim to the balanced wandering of E3 but differs in that it is a reward bonus rather than a policy definition (since we do not have control over the measures) and that it is maintained throughout the entire optimization process as we are interested in diversity.
3 Wasserstein Quality Diversity Imitation Learning (WQDIL)
@@
This section introduces our proposed framework-WQDIL. We first begin with two solutions for the behavior overfitting problem which are called Single-Step Archive Exploration Bonus and Measure Conditioning (Section 3.1). We then introduce Wasserstein Auto-encoders (Section 3.2), and propose a WQDIL instance WAE-WGAIL that applies latent Wasserstein adversarial training for WAE optimization (Section 3.3).
3.1 Mitigating behavior overfitting
@@
Compared to the vast behavior space, expert behaviors tend to be limited in diversity. This poses a problem for learning diverse behaviors, as the reward model in the traditional IL paradigm focuses solely on shaping rewards around specific expert behaviors while neglecting other behavior patterns. We refer to this type of reward as a “behavior-overfitted reward.” Consequently, this issue significantly inhibits the learning of diverse behaviors, since any RL method inherently converges to high-reward behaviors. To address this problem, we introduce a single-step archive-exploration reward bonus and measure-conditioning, which together ensure exploration across the behavior space as well as the sensitivity of the reward function to the local measure information.
@@
In both the bonus and the measure conditioning, we are primarily interested in single-step measures building on the Markovian Measure Proxy mentioned in Section 2.2 to make the measure information more fine-grained. This is not merely for convenience purposes but also as this will yield richer data from the demonstrations, which would otherwise be sparse (e.g. only a few episodic measures since we focus on limited-expert-demonstrations in our setting). Meanwhile, Focusing on single step measure will allow exploring behaviours with similar episodic measure but different single-step measure for potentially yielding higher fitness within a particular behavioral region. This is due to the episodic measure is the average (or sum) of each single-step measure and is less fine-grained.
For the exploration bonus, we establish the single-step archive $\mathcal{A}_{single}$, which corresponds to the state-dependent measure $\delta(s)$. Similar to the behavior archive $\mathcal{A}$, we partition $\mathcal{A}_{single}$ into numerous cells for discretization. Notably, instead of merely recording whether a cell is occupied, we track the visitation count $n_{i}$ for each cell $C_{i}$ in $\mathcal{A}_{single}$. The exploration reward bonus is defined as:
@@
The exploration bonus assigns higher rewards to regions in $\mathcal{A}_{single}$ that are less frequently visited, thereby promoting the agent to explore unseen behavior patterns. Additionally, once a region within the single-step behavior space has been sufficiently explored, the bonus decreases, facilitating the exploitation of that region to discover high-performing policies. However, note that the bonus is defined relative to the exploration of other measures such that the bonus never shrinks to zero for a particular measure. The form in Eq. 3 also avoids extreme values due to the +1 offset in the denominator. With these features together, the reward bonus can effectively mitigate the “behavior-overfitted reward’ issue by always encouraging new behavior patterns, thus facilitating diverse behaviors.
In addition to the exploration bonus, the reward model is also conditioned on the measure, according to $\mathcal{R}(s,a,\delta(s))$. This contributes to a solution to behavior overfitting: since the goal is to form behaviorally diverse policies, different behaviors require different policies to be imitated – and hence correspond to a different reward function.
3.2 Wasserstein auto-encoders
The Wasserstein distance is a formulation that occurs within the context of optimal transport [47]. Kantorovich’s formulation of the problem is given by
@@ 3.2 Wasserstein auto-encoders
Auto-encoders aim to represent input data in a low-dimensional latent space. Modern generative models like variational auto-encoders (VAEs) [26] and generative adversarial networks (GANs) [19] do so by minimizing a discrepancy measure between the data distribution $P_{x}$ and the generative model $P_{G}$. To formulate a Wasserstein based generative model, the form in Eq. 5 is intractable; however, following [7], one can equate the Wasserstein distance to the following WAE objective [46]:
$D_{\mathrm{WAE}}(P_{x},P_{G}):=$ $\inf_{Q(z|x)\in\mathcal{Q}}\mathbb{E}_{P_{x}}\mathbb{E}_{Q(z|x)}\bigl{[}f_{c}\bigl{(}x,G(z)\bigr{)}\bigr{]}$ (6)
@@
WAE-GAN [46] is an instance of WAE when choosing $D_{z}(Q_{z},P_{z})=D_{\mathrm{JS}}(Q_{z},P_{z})$ and using adversarial training to estimate $D_{z}$. Specifically, WAE-GAN uses an adversary and discriminator in the latent space $\mathcal{Z}$ trying to separate “true” points sampled from $P_{z}$ and “fake” ones sampled from $Q_{z}$ [19]. In the imitation learning setting, $P_{z}$ corresponds to the distribution of latent data obtained from the encoded demonstrations while $Q_{z}$ corresponds to the distribution of latent data obtained from the encoded trajectory data from the policy.
Analogously, we propose WAE-WGAN, which is equivalent to WAE-GAN except that it sets the divergence measure to the $1$-Wasserstein distance, i.e. $D_{z}(Q_{z},P_{z})=\mathcal{W}_{1}(Q_{z},P_{z})$. We choose this option based on results on the improved stability during adversarial training [2]. Due to strong duality with $p=1$, the Kantorovic-Rubinstein duality can be applied, i.e.
@@
3.3 WQDIL: WAE Meets Adversarial QDIL
This section introduces our Wasserstein Quality Diversity Imitation Learning (WQDIL) framework. We illustrate the processes of WQDIL in Algorithm 1, which is built on top of the Proximal Policy Gradient Arborescence (PPGA) algorithm [5]. The blue color indicates the WQDIL’s operations that are distinct from the PPGA algorithm.
@@ 3.3 WQDIL: WAE Meets Adversarial QDIL
4:$s_{1},a_{1},\delta(s_{1}),s_{2},a_{2},\delta(s_{2})\dots,s_{k},a_{k},\delta(s_{k})\leftarrow episode$
5:$r_{1},r_{2},\dots,r_{k}\leftarrow\mathcal{R}(\mathbf{s},\mathbf{a},\mathbf{\delta(s)})$$\triangleright$ batch reward
6: compute exploration bonuses $r_{1,exp},r_{2,exp},\dots,r_{k,exp}$ based on equation (3) and (4)
7:For $i=1\rightarrow k$
@@
In WQDIL, we input the initial reward model $\mathcal{R}$ and the expert demonstrations $\mathcal{D}$. In each iteration, we first approximate the gradients $\nabla f$ and $\nabla m$ (in step 4) using the rewards estimated from $\mathcal{R}$ (Algorithm 2). Steps 5-15 are QD optimization options used in PPGA. In step 16, we update the search policy using the reward model $\mathcal{R}$. Step 17 is a key step, in which we update the reward model $\mathcal{R}$ using the input expert demonstrations $\mathcal{D}$. We have explored a few algorithms in the following for updating the reward model $\mathcal{R}$ in WQDIL.
@@
WAE-WGAIL. Although WAE-GAIL achieves better stability and higher QD performance than GAIL in some tasks, the improvements are not consistent. To further improve WAE-GAIL, we propose to apply Wasserstein adversarial training in the latent space of WAE, which is analogous to the WAE-WGAN proposed in Section 3.2. The loss function of the auto-encoder in WAE-WGAIL is given by
@@
$+\lambda\mathbb{E}_{x^{e}\sim\mathcal{D},x^{\pi}\sim\pi}\left[D_{\mathcal{W}}(G_{\psi}(x^{e}))-D_{\mathcal{W}}(G_{\psi}(x^{\pi}))\right],$ (8)
where $\phi$ and $\psi$ are the parameters of the encoder and decoder, respectively, $x=(s,a)$ is the state-action pair, $Q_{\phi}$ is the encoder, $G_{\psi}$ is the decoder, and $D_{\mathcal{W}}$ is the Wasserstein discriminator. The resulting algorithm is called WAE-WGAIL. Algorithm 6 in Appendix A summarizes the details of WAE-WGAIL.
Algorithm 3 mCWAE-WGAIL.
1:Input: Expert emonstrations $\mathcal{D}$, Current policy $\pi_{\theta}$, Regularization coefficient $\lambda>0$.
2:Initialize the parameters of the encoder $Q_{\phi}$,
4:Method: Update reward model in Algorithm 1 (step 17)
5:def update($\mathcal{D},\pi_{\theta}$):
10: Update $D_{\mathcal{W}}$ by ascending:
11: Update $Q_{\phi}$ and $G_{\psi}$ by descending:
12: Repeat until the model converges or the number of epochs is reached
13:return Updated $Q_{\phi}$, $G_{\psi}$ and $D_{\mathcal{W}}$
@@
4 Experiments
4.1 Experiment Setup
We evaluate our framework on three popular MuJoCo [45] environments: Halfcheetah, Humanoid, and Walker2d. The measure function is the number of times a leg contacts with the ground, divided by the trajectory length. We implement all the methods based on the PPGA implementation, which utilizes the Brax simulator [17] with QDax wrappers for measure calculation [27]. All Experiments were conducted on A40 GPUs for three runs per task.
4.2 Demonstrations
We use a policy archive obtained by PPGA to generate expert demonstrations. To follow real-world scenario with limited demonstrations, we first sample the top 500 high-performance elites from the archive as candidate pool. Then from this pool, we select a few demonstrations such that they are as diverse as possible. This process results in 4 diverse demonstrations (episodes) per environment. Figure 3 visualizes the demonstrators in the policy archive, and Table 6 in Appendix C provides the statistical properties of the demonstrations.
@@ 4.2 Demonstrations
Figure 3: Visualization of the behavior space. Green indicates the full expert behavior space, blue indicates the selected top-500 elites, and red indicates the selected demonstrators.
4.3 Performance
Figure 4: Visualization of the policy archive of Expert, GAIL, WAE-GAIL, WAE-WGAIL, WAE-WGAIL-Bonus and mCWAE-WGAIL-Bonus on Humanoid. The color indicates the cumulative rewards of best performing policy in the archive cells.
Figure 5: Learning curve comparison of our mCWAE-WGAIL-Bonus against WAE-WGAIL and state-of-the-art IL methods.
To validate the effectiveness of WQIL with Single-Step Archive Exploration, we compare our mCVAE-WAIL-Bonus against state-of-the-art IL methods within QDIL framework. These methods includes our WAE-WGAIL (Algorithm 6, Appendix A), GAIL [23], PWIL [12], AIRL [18], MaxEntIRL [52] and GIRIL [49]. We apply a gradient penalty to all the methods that require the online update of the reward model. Appendix B summarizes the hyperparameters.
We compare the performance under four QD metrics [5]: 1) QD-Score, the sum of scores of all nonempty cells in the archive. QD-score is the most important metric in QDIL as it aligns with the objective of QDIL as in equation (2); 2) Coverage, the percentage of nonempty cells, indicating the algorithm’s ability to discover diverse behaviors; 3) Best Reward, the highest score found by the algorithm; and 4) Average Reward, the mean score of all nonempty cells, reflecting the ability to discover both diverse and high-performing policies. We use the true reward functions to calculate these metrics.

### v2

@@ 2.5. Encouraging exploration in RL and IL
Exploration bonuses have a long history in bandits and reinforcement learning as part of the “optimism in face of uncertainty” principle, and include but are not limited to visitation count based approaches. In bandits, upper confidence bound (UCB, e.g. Auer et al. (2002)) computes a bonus based on $\sqrt{\frac{2\ln(t)}{n(a_{t})}}$ for action $a_{t}$ at time $t$. In tabular reinforcement learning, upper confidence reinforcement learning (UCRL) techniques (e.g. Jaksch et al. (2010)) provide upper confidence bounds to the reward based on a bonus proportional to $\sqrt{\frac{1}{n(s_{t},a_{t})}}$. R-MAX Brafman and Tennenholtz (2002) and E3 Kearns and Singh (2002) explicitly distinguish between known and unknown states, where in unknown states R-MAX assigns the maximal reward of the MDP and E3 performs balanced wandering (i.e. taking the action with lowest visitation count for the current state). Related concepts have also been applied in deep reinforcement learning Tang et al. (2017); Henaff et al. (2022); Machado et al. (2020); Seo et al. (2021). Related to exploration bonuses, recent techniques in imitation learning also considered curiosity-based exploration (Pathak et al., 2017; Burda et al., 2019; Yu et al., 2020), which relates to discrepancy between the transition model’s prediction and the true transition. For instance, GIRIL (Yu et al., 2020) computes rewards offline by pretraining a reward model using a conditional VAE (Sohn et al., 2015).
@@
Our measure bonus differs from these approaches in that it is designed to explore the behavior space rather than the (state-)action space. Further, contrasting to UCRL type bonuses, due to dividing by the proportion of visits, rather than the count, our measure bonus does not shrink to zero such that underexplored behaviors will continue to receive a bonus. This is somewhat comparable in aim to the balanced wandering of E3 but differs in that it is a reward bonus rather than a policy definition (since we do not have control over the measures) and that it is maintained throughout the entire optimization process as we are interested in behavioral diversity.
3. Wasserstein Quality Diversity Imitation Learning (WQDIL)
@@
This section introduces our proposed framework-WQDIL. We first introduce the behavior overfitted reward problem and propose two solutions called Single-Step Archive Exploration Bonus and Measure Conditioning (Section 3.1). We then introduce Wasserstein Auto-encoders (Section 3.2), and propose a WQDIL instance WAE-WGAIL that applies latent Wasserstein adversarial training for WAE optimization (Section 3.3).
3.1. Mitigating behavior overfitting
@@
Compared to the vast behavior space, expert behaviors tend to be limited in diversity. This poses a problem for learning diverse behaviors, as the reward model in the traditional IL paradigm focuses solely on shaping rewards around specific expert behaviors while neglecting other behavior patterns. This presents a challenge for balancing exploration and exploitation and makes reward functions local in that they only count a single behavior as high-performing Wan et al. (2024). We refer to this as the “behavior-overfitted reward” problem. This problem significantly inhibits the learning of diverse behaviors, since any RL method inherently converges to high-reward behaviors. To address this problem, we introduce a single-step archive-exploration reward bonus and measure-conditioning, which together ensure exploration across the behavior space as well as the sensitivity of the reward function to the local measure information.
@@
In both the bonus and the measure conditioning, we are primarily interested in single-step measures building on the Markovian Measure Proxy mentioned in Section 2.2 to make the measure information more fine-grained. This is not merely for convenience purposes but also as this will yield richer data from the demonstrations, which would otherwise be sparse (e.g. only a few episodic measures since we focus on limited-expert-demonstrations in our setting). Meanwhile, focusing on single step measure will allow exploring behaviors with similar episodic measure but different single-step measure for potentially yielding higher fitness within a particular behavioral region. This is because the episodic measure is the average (or sum) of each single-step measure and is less fine-grained.
For the exploration bonus, we establish the single-step archive $\mathcal{A}_{single}$, which corresponds to the state-dependent measure $\delta(s)$. Similar to the behavior archive $\mathcal{A}$, we partition $\mathcal{A}_{single}$ into numerous cells for discretization. Notably, instead of merely recording whether a cell is occupied, we track the visitation count $n_{i}$ for each cell $C_{i}$ in $\mathcal{A}_{single}$. The exploration reward bonus is formulated as
@@
The exploration bonus assigns higher rewards to regions in $\mathcal{A}_{single}$ that are less frequently visited, thereby promoting the agent to explore unseen behavior patterns. Additionally, once a region within the single-step behavior space has been sufficiently explored, the bonus decreases, facilitating the exploitation of that region to discover high-performing policies. However, note that the bonus is defined relative to the exploration of other measures such that the bonus never shrinks to zero for a particular measure. The form in Eq. 3 also avoids extreme values due to the +1 offset in the denominator. With these features together, the reward bonus can effectively mitigate the “behavior-overfitted reward” issue by always encouraging new behavior patterns, thus facilitating diverse behaviors.
In addition to the exploration bonus, the reward model is also conditioned on the measure. Specifically, using GAIL for imitation learning, the reward model is formulated as
~$\tilde{\mathcal{R}}(s,a,\delta(s))$ $=\max_{\pi}\min_{D}\mathbb{E}_{(s,a)\sim\mathcal{D}}[-\log D(s,a,\delta(s))]$
$+\mathbb{E}_{(s,a)\sim\pi}[-\log(1-D(s,a,\delta(s)))].$ (5)
This contributes to a solution to behavior overfitting: since the goal is to form behaviorally diverse policies, different behaviors require different policies to be imitated – and hence correspond to a different reward function. Combining with the exploration bonus, we obtain the combined reward model
$\mathcal{R}(s,a,\delta(s))=\tilde{\mathcal{R}}(s,a,\delta(s))+r_{exp}(s,a,\delta(s)).$ (6)
3.2. Wasserstein auto-encoders
The Wasserstein distance is a formulation that occurs within the context of optimal transport Villani (2003). Kantorovich’s formulation of the problem is given by
@@ 3.2. Wasserstein auto-encoders
Auto-encoders aim to represent input data in a low-dimensional latent space. Modern generative models like variational auto-encoders (VAEs) Kingma and Welling (2014) and generative adversarial networks (GANs) Goodfellow et al. (2014) do so by minimizing a discrepancy measure between the data distribution $P_{x}$ and the generative model $P_{G}$. To formulate a Wasserstein based generative model, the form in Eq. 7 is intractable; however, following Bousquet et al. (2017), one can equate the Wasserstein distance to the following WAE objective Tolstikhin et al. (2018):
$D_{\mathrm{WAE}}(P_{x},P_{G}):=$ $\inf_{Q(z|x)\in\mathcal{Q}}\mathbb{E}_{P_{x}}\mathbb{E}_{Q(z|x)}\bigl{[}f_{c}\bigl{(}x,P_{G}(z)\bigr{)}\bigr{]}$ (8)
@@
WAE-GAN (Tolstikhin et al., 2018) is an instance of WAE when choosing $D_{z}(Q_{z},P_{z})=D_{\mathrm{JS}}(Q_{z},P_{z})$ and using adversarial training to estimate $D_{z}$. Specifically, WAE-GAN uses an adversary and discriminator in the latent space $\mathcal{Z}$ trying to separate “true” points sampled from $P_{z}$ and “fake” ones sampled from $Q_{z}$ Goodfellow et al. (2014). In the imitation learning setting, $P_{z}$ corresponds to the distribution of latent data obtained from the encoded demonstrations while $Q_{z}$ corresponds to the distribution of latent data obtained from the encoded trajectory data from the policy.
It has been noted previously that WAE-GAN still suffers from instabilities due to the mode collapse property of GANs Kolouri et al. (2019). Therefore, we propose WAE-WGAN, which is equivalent to WAE-GAN except that it sets the divergence measure to the $1$-Wasserstein distance, i.e. $D_{z}(Q_{z},P_{z})=\mathcal{W}_{1}(Q_{z},P_{z})$. We choose this option based on results on the improved stability during adversarial training Arjovsky et al. (2017). Due to strong duality with $p=1$, the Kantorovic-Rubinstein duality can be applied, i.e.
@@
3.3. WQDIL: WAE Meets Adversarial QDIL
This section introduces our Wasserstein Quality Diversity Imitation Learning (WQDIL) framework, which is built on top of the Proximal Policy Gradient Arborescence (PPGA) algorithm (Batra et al., 2023b). The pseudo-code of WQDIL is given in Algorithm 1, where the blue color indicates the WQDIL’s operations that are distinct from the PPGA algorithm.
@@ 3.3. WQDIL: WAE Meets Adversarial QDIL
4:$s_{1},a_{1},\delta(s_{1}),s_{2},a_{2},\delta(s_{2})\dots,s_{k},a_{k},\delta(s_{k})\leftarrow\texttt{episode}()$
5:$r_{1},r_{2},\dots,r_{k}\leftarrow\tilde{\mathcal{R}}(s,a,\delta(s))$triangleright batch reward (see Eq.3.1)
6: compute exploration bonuses $r_{1,exp},r_{2,exp},\dots,r_{k,exp}$triangleright (see Eq.3)
7:For $i=1,\dots,k$
@@
In WQDIL, we first initialize the reward model $\mathcal{R}$ and input the expert demonstrations $\mathcal{D}$. Then we repeat iterations as follows. First, approximate the gradients $\nabla f$ and $\nabla m$ (in l.4) using the rewards estimated from $\mathcal{R}$ (Algorithm 2). Lines 5-15 are QD optimization options used in PPGA. In line 16, we update the search policy using the reward model $\mathcal{R}$. Line 17 updates the reward model $\mathcal{R}$ using the input expert demonstrations $\mathcal{D}$. We now discuss different algorithms for updating the reward model $\mathcal{R}$ in WQDIL.
@@
WAE-WGAIL. Although WAE-GAIL achieves better stability and higher QD performance than GAIL in some tasks, the improvements are not consistent. To further improve WAE-GAIL, we propose to apply Wasserstein adversarial training in the latent space of WAE, which is analogous to the WAE-WGAN proposed in Section 3.2. The loss function of the auto-encoder in WAE-WGAIL is based on the reconstruction loss for expert demonstrations, the reconstruction loss for policy data, and the Wasserstein distance between latent variables of expert demonstrations and policy data:
@@
$+\lambda\mathbb{E}_{z^{e},z^{\pi}}\left[D_{\mathcal{W}}(z^{e})-D_{\mathcal{W}}(z^{\pi})\right],$ (10)
where $\phi$ and $\psi$ are the parameters of the encoder and decoder, respectively, $x=(s,a)$ is the state-action pair, $Q_{\phi}$ is the encoder, $G_{\psi}$ is the decoder, and $D_{\mathcal{W}}$ is the Wasserstein discriminator. Algorithm 5 in Appendix A summarizes the details of WAE-WGAIL.
@@
Figure 3. Visualization of the demonstrations obtained from PPGA archives. The x and y axes are the proportions of time leg 1 and 2, respectively, touch the ground. Green indicates the full expert behavior space, blue indicates the selected top-500 elites, and red indicates the selected demonstrators.
Figure 4. Visualization of the policy archive of Expert, GAIL, WAE-GAIL, WAE-WGAIL, WAE-WGAIL-Bonus and mCWAE-WGAIL-Bonus on Humanoid. The color indicates the cumulative rewards of best performing policy in the archive cells.
Figure 5. Learning curve comparison of our mCWAE-WGAIL-Bonus against WAE-WGAIL and state-of-the-art IL methods. The curves (and shaded areas) represent the means (and standard deviations) of the algorithms. Columns indicate the metric (QD-Score, Coverage, Best Reward, and Average Reward) while the rows represent the different benchmarks (Halfcheetah, Walker2d, and Humanoid).
4. Experiments
4.1. Experiment Setup
We evaluate our framework on three popular MuJoCo (Todorov et al., 2012) environments: Halfcheetah, Humanoid, and Walker2d. The measure function is the number of times a leg contacts with the ground, divided by the trajectory length. We implement all the methods based on the PPGA implementation, which utilizes the Brax simulator (Freeman et al., 2021) with QDax wrappers for measure calculation (Lim et al., 2022). All experiments were conducted on A40 GPUs for three runs per task.
4.2. Demonstrations
We use a policy archive obtained by PPGA to generate expert demonstrations. To follow real-world scenario with limited demonstrations, we first sample the top 500 high-performance elites from the archive as candidate pool and then select 4 policies from these candidates such that they are as diverse as possible. We generate 1 demonstration for each selected policy, resulting in the final 4 demonstrations for QDIL. This process results in 4 diverse demonstrations (episodes) per environment. Figure 3 visualizes the demonstrators in the policy archive, and Table 4 in Appendix C provides the statistical properties of the demonstrations.
4.3. Performance
To validate the effectiveness of WQIL with Single-Step Archive Exploration, we compare our mCVAE-WAIL-Bonus against state-of-the-art IL methods within QDIL framework. These methods includes our WAE-WGAIL (Algorithm 5, Appendix A), GAIL Ho and Ermon (2016), PWIL Dadashi et al. (2021), AIRL Fu et al. (2017), MaxEntIRL Ziebart et al. (2008) and GIRIL Yu et al. (2020). We apply a gradient penalty to all the methods that require the online update of the reward model. Appendix B summarizes the hyperparameters.
We compare the performance under four QD metrics Batra et al. (2023b): 1) QD-Score, the sum of scores of all nonempty cells in the archive. QD-score is the most important metric in QDIL as it aligns with the objective of QDIL as in equation (2); 2) Coverage, the percentage of nonempty cells, indicating the algorithm’s ability to discover diverse behaviors; 3) Best Reward, the highest score found by the algorithm; and 4) Average Reward, the mean score of all nonempty cells, reflecting the ability to discover both diverse and high-performing policies. We use the true reward functions to calculate these metrics.